OpenStudy (anonymous):

Hint: Pay attention to the units of measure. You may have to convert from feet to miles several times in this assignment. You can use 1 mile = 5,280 feet for your conversions. 1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. quation is w=Cr^-2

OpenStudy (anonymous):

u only asked 1 question in ur whole thing. Answer: yes i did know that

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the question is solve the equation for r

OpenStudy (anonymous):

o ok well then w=Cr^(-2) w/C=1/(r^2) C/w=r^2 \[\pm \sqrt{C/w} =r\] radius can only be positive so final answer is \[r=\sqrt{C/w}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

was there more to the question? like solve for a value?

OpenStudy (anonymous):

i am posting it in parts it is too big to put all together